# Conceptualization of this study, Methodology, Software

### v1

@@ Abstract
Continual learning (CL) refers to the ability of an algorithm to continuously and incrementally acquire new knowledge from its environment while retaining previously learned information. A model trained on one data modality often fails when tested with a different modality. A straightforward approach might be to fuse the two modalities by concatenating their features and training the model on the fused data. However, this requires retraining the model from scratch each time it encounters a new domain. In this paper, we introduce a continual learning algorithm capable of incrementally learning different data modalities by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce, yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful. The proposed algorithm is efficient and only requires storing prototypes for each class. We evaluate the algorithm’s effectiveness on a challenging custom multimodal dataset comprising of tactile data from a soft pneumatic gripper, and visual data from non-stationary images of objects extracted from video sequences. Additionally, we conduct an ablation study on the custom dataset and the Core50 dataset to highlight the contributions of different components of the algorithm. To further demonstrate the robustness of the algorithm, we perform a real-time experiment for object classification using the soft gripper and an external independent camera setup, all synchronized with the Robot Operating System (ROS) framework.
@@
[…]
@@ 1 Introduction
In recent years, machine learning models have matched or even exceeded human-level performance on tasks such as image classification Ahmad et al., object recognition Zou et al., natural language processing (NLP) Otter et al., playing games in simulation, etc. While these models excel in static tasks where the data distribution remains constant, they encounter difficulties in adapting to changing environments and need to restart the training process whenever new data is introduced. On the other hand, humans are adept at learning from dynamic environments, continuously acquiring, updating, and applying knowledge over time. We expect AI systems to adapt in a similar fashion.
Continual learning (CL) Wang et al., Lesort et al., also known as lifelong learning or incremental learning, is a machine learning subfield that focuses on progressively training models on a continuous stream of data to accumulate and retain knowledge over time. However, a significant obstacle for these artificial models is their tendency to forget previously acquired skills when exposed to new tasks or data distributions. This issue, known as catastrophic forgetting, often leads to a substantial drop in performance as the new information overwrites the old knowledge or alters the data distribution. A naive solution to this problem is to save all data, shuffle it, and retrain the model offline from scratch. While this method effectively addresses catastrophic forgetting, it is time-inefficient, suboptimal, and unsustainable for modern large-scale models that require significant computational power. Continual learning aims to develop methods that balance performance and efficiency more effectively. Robots operating in unstructured real world environments will encounter new tasks and challenges over time, necessitating capabilities that cannot be fully anticipated from the outset. These robots need to learn continually, acquiring new skills without forgetting previously learned ones.
@@
A model trained in one domain will be ineffective when deployed in a completely different domain or used with a different data modality than it was originally trained on. One straightforward approach might be to use separate models for each data modality or to jointly train a single model on all modalities offline. However, this is not a feasible solution for agents with limited memory and computational resources that learns new domains through active exploration and interaction with its environment. Another challenge is the limited availability of labeled data for supervised incremental training. A large research effort is dedicated towards learning from unlabeled data, which is often more readily available than labeled data in practical applications. Within this effort, the field of semi-supervised learning (SSL) Van Engelen and Hoos has recently demonstrated the most promising results. In this setting, an agent can utilize a significant amount of unlabeled data from its environment. It can employ the knowledge gained from labeled data to learn new objects of the same class or to enrich its understanding of previously learned classes using the unlabeled data.
@@
In this paper, we consider a realistic continual learning problem where a single continual learning model learns different modalities of data incrementally by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful. We mainly experiment with two different SSL conditions: Radom images, where for each object of the same parent class we randomly sample some images from the dataset that are not included in the training data; Unique objects, where we sample the images of new unseen objects that belong to the same parent class. More details about the SSL conditions is provided in section 3.3.
@@
To perform a multimodal CL experiment, we leverage two different modalities of data mainly the tactile data coming from sensors integrated with a soft gripper and the visual data from an independent camera setup. The CL model is designed to learn the classes in both the domains incrementally. To address this issue of multimodal CL, we build on the feature covariance-aware metric (FeCAM) algorithm by Goswami et al. and propose an extended version of the algorithm that employs an incremental online semi-supervised learning process for each task. We call it exFeCAM or extended FeCAM algorithm. Additionally, our algorithm includes an intra-layer feature representation mechanism to learn more generalized feature maps for each class. To the best of our knowledge, we are the first to use a continual learning model in a multimodal data setting where an agent learns both tactile data and vision data provided by respective sensors. The main contributions of this article are summarized as follows:
@@
    We present an online non-exemplar continual learning algorithm with SSL capabilities. In addition, to enhance the feature maps from the pre-trained layers, we introduce intra-layer feature representations.
@@
    We present a new multimodal non-iid dataset for real world continual learning applications.
@@
    We evaluate our algorithm on the custom multimodal dataset as well as the Core50 dataset Lomonaco and Maltoni and perform an ablation study to verify the effectiveness of the various parts of the algorithm.
@@
    We also perform a real-time experiment for object classification on the soft pneumatic gripper equipped with sensors and a separate camera setup, all synchronised using the ROS framework Koubaa et al.. More details about the experiment is provided in section 6.
The paper is structured as follows: In Section 2, an overview of the related works in CL, SSL, and multimodality is presented, with a focus on the recent advancements in the literature. Section 3 provides a detailed description of the experimental setup that has been adopted discussing the different cl strategies, data accumulation steps, and evaluation metrics. Following this, the proposed methodology is presented in Section 4, which includes the model architecture and the framework used in this study. The evaluation of the proposed methodology on custom multimodality dataset, as well as the Core50 dataset is provided in section 5. Moreover, in section 6 the paper showcases the real-time application of the proposed algorithm. Finally, a summary noting the advantages and disadvantages of the proposed method along with the scope for improvements and possible developments in the future are stated in the last section.
@@ 2.1 Continual Learning Strategies
Continual learning strategies can be broadly categorized into four major types: architectural, regularization, rehearsal, and hybrid strategies, each with its own advantages and disadvantages. Architectural strategy either employs a dynamic modular network where new layers gets added upon encounter of new tasks Rusu et al., Fernando et al., Rajasegaran et al., Aljundi et al. or performs parameter isolation to allocate a dedicated separate sub-space for each task within the same network Mallya et al., Serra et al., Mallya and Lazebnik, Ahn et al.. The simplest form of architectural regularization involves freezing certain weights in the network to ensure they remain unchanged during training on new tasks. A major drawback of the architectural strategy is the increase in the size of the architecture as new tasks are learned or the potential saturation of the architecture if a fixed model is used.
@@
Regularization strategies can be grouped into two: weight regularization and function regularization. Weight regularization constrains the update of the network weights by adding explicit regularization terms to the loss function, penalizing changes in each parameter based on its importance Kirkpatrick et al., Zenke et al., Chaudhry et al., Aljundi et al.. Function regularization, on the other hand, employs distillation methods to maintain the knowledge between the various layers of the network. This strategy typically employs the previously-learned model as the teacher and the currently model as the student, while implementing knowledge distillation (KD) techniques Hinton et al., Li and Hoiem, Rebuffi et al. Dhar et al., Lee et al.. However, regularising the model can make it more challenging for the network to acquire new knowledge.
@@
Rehearsal Lopez-Paz and Ranzato, Bang et al., Kumari et al., Ebrahimi et al., Saha and Roy & pseudo-rehearsal, Shin et al., Wu et al., Ayub and Wagner, Pfülb and Gepperth, Wang et al. approaches offer the most simple yet effective approach by storing a subset of the previously learned data in a memory buffer or by training a generative model to account for the previous information and replaying it during training on new tasks. However, storing the images in a replay buffer for rehearsal can raise privacy concerns. On the other hand, while using a generative model somewhat addresses the privacy issues but may result in increased training time and computational demands.
@@
Hybrid strategies Van de Ven et al., Liu et al., Ye and Bors are a combination of any of the other three strategies to compensate for catastrophic forgetting. SynapNet Kushawaha et al. uses a triple model architecture incorporating knowledge distillation along with a VAE-based pseudo-episodic memory for rehearsal and a sleep phase for memory re-organization. DualNet Pham et al. uses a fast learner for supervised learning and a slow learner for unsupervised learning of task-agnostic general representation using semi-supervised learning. ICARL Rebuffi et al. includes an external fixed memory to store a subset of old task data based on an elaborated sample selection procedure and then employs a distillation step acting as a regularization.
@@
Non-exemplar continual learning strategies Goswami et al., Petit et al., Smith et al., Hayes et al. use the idea of replay-free sequential training, instead of using a modular network or regularizing the optimization function. Exemplar-free approaches store the prototypes of the past and new classes to compensate for catastrophic forgetting of the previously acquired knowledge. A pre-trained feature extractor provides the feature maps which are then used to calculate the mean and the covariance matrix for the respective classes.
@@ 2.2 Semi-Supervised Learning
Semi-supervised learning describes a class of algorithms designed to learn from both labeled and unlabeled data, typically assumed to be sampled from the same or similar distributions Yan et al., Mawuli et al.. One of the critical challenges in semi-supervised learning is how to effectively utilize labeled and unlabeled data while avoiding overfitting, preventing the effects of noise on the model, and ensuring effective model performance enhancement. Initial breakthroughs in semi-supervised learning with deep neural networks leveraged generative models such as autoencoders Rasmus et al., variational autoencoders Kingma et al., and generative adversarial networks Odena.
Some approaches in semi-supervised learning Oliver et al., Berthelot et al., Lee et al., Kingma et al., Kuo et al. also involve balancing a supervised loss ($l_{s}$) applied to the labeled data and an unsupervised consistency regularization loss $l_{ul}$ computed on the unlabeled data. Consistency regularization aims to improve model resilience and maintain label distribution, even in the presence of noisy images. It measures the discrepancy between predictions made on perturbed unlabeled data points. Approaches of this kind include the $\pi$-model Laine and Aila which augments the input data with image-based noise and self-regularizes through an additional consistency loss, mean teacher Tarvainen and Valpola, which employs an exponential moving average of parameters to regularize the model. Recently, fast-SWA Athiwaratkun et al. showed improved results by training with cyclic learning rates and measuring discrepancy with an ensemble of predictions from multiple checkpoints.
Two additional important approaches for semi-supervised learning, which have shown success both in the context of deep neural networks and other types of models are: self-training and conditional entropy minimization. Self-training also known as pseudo-labeling Lee et al. involves assigning classes to unlabeled data by making predictions from a model trained only on labeled data. Furthermore, co-training Xie et al. attempts to analyze data from different perspectives to solve the accumulative mistake issue during self-training. On the other hand, in case of conditional entropy minimization Grandvalet and Bengio, all unlabeled examples are encouraged to make confident predictions on some class. This doesn’t necessarily mean the predictions are correct (because the true labels for the unlabeled data are unknown), but the model is incentivized to be decisive rather than uncertain. In this paper, we use the concept of cosine similarity between labeled and unlabeled feature representations for pseudo-labeling of the unsupervised data. We refer to Van Engelen and Hoos for a comprehensive review on the topic.
@@ 2.3 Multimodality
Our experience of the world is inherently multimodal we perceive objects through vision, hear sounds, feel textures, smell odors, and taste flavors. To develop an AI agent with similar capabilities, researchers have explored various methods Donato et al., Babadian et al..
@@
The simplest method to compensate multiple domains is using the concept of joint representation where the unimodal signals are combined in the same representation space by simple concatenation Baltrušaitis et al.. Ngiam et al. used autoencoders for multimodal domains. They employed stacked denoising autoencoders to represent each modality individually and subsequently fusing them into a unified multimodal representation using an additional autoencoder layer.
Similarly, Silberer and Lapata suggested a multimodal autoencoder for semantic concept grounding. They incorporated a reconstruction loss for training the representation and added a term in the loss function to predict object labels using the representation. Srivastava and Salakhutdinov introduced multimodal deep belief networks (DBNs) and multimodal deep boltzmann machines (DBMs) Srivastava and Salakhutdinov as multimodal representations. Kim et al. used a deep belief network for each modality and then combined them into joint representation for audiovisual emotion recognition. Huang and Kingsbury used a similar model for audio-visual speech recognition (AVSR), and Wu and Shao used it for audio and skeleton joint based gesture recognition. Ouyang et al. explored the use of multimodal DBMs for the task of human pose estimation from multi-view data. They demonstrated that integrating the data at a later stage after unimodal data, underwent nonlinear transformations, which was beneficial for the model. An alternative to a joint multimodal representation is a coordinated representation where instead of projecting the modalities together into a joint space, separate latent representations are learned for each modality and are coordinated through a constraint Frome et al..
Sarfraz et al. recently utilized a multimodal CL dataset to learn complementary information from two different modalities. They employed the VGGSound dataset, which includes 10-second video clips along with their corresponding audio. On the other hand, Xu et al. developed an egocentric multimodal dataset using smart glasses equipped with sensors, combining data from cameras, accelerometers, and gyroscopes. This dataset covers 32 types of daily activities performed by 10 participants. They utilized the temporal binding network (TBN) Kazakos et al. to fuse the information from different sensors into single feature vector for further training and feature extraction.
One of the major drawbacks of these methods is their inability to incrementally learn new domains. For example, if we train our model on two domains (sensor signals and images) and use a joint representation to learn these two modalities, introducing a new data modality later would necessitate re-training the model from scratch. Conversely, with our proposed exFeCAM model, we can add the new modality as a different domain and incrementally train the model without starting the training process anew. Additionally, our algorithm eliminates the necessity of storing data in memory for rehearsal, effectively addressing privacy concerns and storage challenges commonly encountered in exemplar-based CL methods.
With recent advancements in large language models (LLMs) and the introduction of GPT-4 Achiam et al. and its variants, it is now possible to utilize the data coming from vision, audio, and text signals into a single model. However, these models face significant challenges, including high computational and memory demands and a lack of continual learning capabilities.
@@
Figure 1: Multimodal data comprising of two different domains, domain 1 contains the sensor information for different objects and domain 2 contains the images of those objects in different orientations, lightning conditions and different level of occlusion by the operator. Moving from left to right the model trains in a class-incremental manner, while going from top-bottom it trains in a domain-incremental manner.

### v2

# Continual Learning for Multimodal Data Fusion of a Soft Gripper

@@ Abstract
Continual learning (CL) refers to the ability of an algorithm to continuously and incrementally acquire new knowledge from its environment while preserving previously learned information. Traditional models trained on a single data modality often fail to generalize when exposed to different modalities. A simple strategy to address this limitation is to fuse multiple modalities by concatenating their features and training the model on the combined representation. However, this generally requires re-training the model from scratch whenever a new domain is introduced. In this paper, we propose a continual learning algorithm that incrementally learns from multiple data modalities by combining class-incremental and domain-incremental learning settings. We validate our algorithm on a custom multimodal dataset composed of tactile signals collected from a soft sensorized pneumatic gripper and visual data consisting of non-stationary object images captured from video sequences. Additionally, we evaluate our method on a subset of the publicly available VGGSound dataset, which integrates both visual and audio signals corresponding to various real-world activities. To further demonstrate the robustness and real-time applicability of our approach, we conduct an object classification experiment using a soft sensorized gripper and an external camera setup, all synchronized through the Robot Operating System (ROS) framework.
@@
Continual Learning, Class prototypes, Multimodality, Robot operating system, Semi-supervised learning, Soft gripper, Tactile data
Nilay Kushawaha* Egidio Falotico
{affiliations}
Nilay Kushawaha
Address[1]: The BioRobotics Institute, Scuola Superiore Sant’Anna, Pontedera 56025, Italy
Address[2]: Department of Excellence in Robotics and AI, Scuola Superiore Sant’Anna, Pisa 56127, Italy Email Address: nilay.kushawaha@santannapisa.it
Egidio Falotico
Address[1]: The BioRobotics Institute, Scuola Superiore Sant’Anna, Pontedera 56025, Italy
Address[2]: Department of Excellence in Robotics and AI, Scuola Superiore Sant’Anna, Pisa 56127, Italy
@@ 1 Introduction
In recent years, machine learning models have matched or even exceeded human-level performance on various tasks, such as image classification [1], object recognition [2], natural language processing (NLP) [3], and simulated game play [4]. Despite their success in static settings where data distributions remain stable, these models struggle to adapt to evolving environments and often require re-training the network from scratch when new data becomes available. In contrast, humans demonstrate remarkable adaptability, continually learning, updating, and applying knowledge over time. Our goal is to develop AI systems that can replicate this level of continual adaptability.
Continual learning (CL) [5], also referred to as lifelong or incremental learning is a subfield of machine learning that focuses on training models progressively on a stream of data, allowing it to accumulate and retain knowledge over time. However, a major challenge for such models is their susceptibility to catastrophic forgetting of the previously acquired skills when learning new tasks/experiences or encountering shifts in data distribution. This often results in significant performance degradation, as new information overwrites existing knowledge. A straightforward but impractical solution is to store all past data, shuffle it, and re-train the model offline from scratch. While this can mitigate forgetting, it is computationally expensive, inefficient, and unsustainable for large-scale models. The goal of continual learning is to develop more scalable and efficient approaches that can maintain performance over time.
@@
A model trained in one domain often performs poorly when deployed in a completely different domain or when applied to a different data modality than it was originally designed for. A common approach to address this issue is to either train separate models for each modality or to jointly train a single model offline on all modalities. However, these strategies are impractical for agents with limited memory and computational resources, especially those that learn new domains through active exploration and interaction with their environment. Another significant challenge is the scarcity of labeled data for supervised incremental learning. In contrast, unlabeled data is typically more abundantly available in real-world applications. To address this imbalance, semi-supervised learning (SSL) [6] has emerged as a promising solution. In this setting, an agent leverages a small set of labeled data to extract useful knowledge, which can then be applied to learn new instances of the same class or to enhance it’s understanding of previously encountered classes using the unlabeled data.
@@
In this paper, we consider a realistic CL problem where a single CL model learns different modalities of data incrementally by leveraging both class-incremental and domain-incremental learning scenarios in an artificial environment where labeled data is scarce yet non-iid (independent and identical distribution) unlabeled data from the environment is plentiful. We also experiment with two different SSL conditions: Radom images, where for each object of the same parent class we randomly sample some images from the dataset that are not included in the training data; Unique objects, where we sample the images of new unseen objects that belong to the same parent class. More details about the SSL conditions is provided in section 3.3.
@@
For our multimodal CL experiment, we employ two distinct types of dataset. The first setup involves tactile input from sensors embedded within a soft gripper, alongside visual data captured by an independent camera system. For the second experiment, we use the VGGSound dataset [7], which offers paired image and audio modalities extracted from video clips. Our CL model is designed to incrementally learn class representations across these different domains. To address the challenges associated with multimodal CL, we build upon the Feature Covariance-Aware Metric (FeCAM) algorithm introduced by [8]. We propose an enhanced version, referred to as exFeCAM (extended FeCAM), which incorporates an online, task-wise semi-supervised learning strategy. Furthermore, our approach introduces an intra-layer feature representation mechanism that enables the learning of more generalized feature maps for each class. To the best of our knowledge, this is the first work to apply continual learning in a multimodal setting, enabling an agent to learn from different modalities in a domain-incremental manner. The key contributions of this work are as follows:
@@
    We present an online non-exemplar CL algorithm with SSL capabilities. In addition, we incorporate intra-layer feature representation technique to obtain generalized feature maps.
@@
    We present a new multimodal non-iid dataset for real-world CL applications.
@@
    We evaluate our algorithm on the custom multimodal dataset as well as the VGGSound dataset [7] and perform an ablation study to verify the effectiveness of the various parts of the algorithm.
@@
    We also perform a real-time experiment for object classification using a soft pneumatic gripper equipped with sensors and a separate camera setup, all synchronized using the ROS framework [9]. More details about the experiment is provided in section 6.
The paper is structured as follows: In Section 2, an overview of the related works in CL, SSL, and multimodality is presented, with a focus on the recent advancements in the literature. Section 3 provides a detailed description of the experimental setup that has been adopted discussing the different CL strategies, data accumulation steps, and evaluation metrics. Following this, the proposed methodology is presented in Section 4, which includes the model architecture and the framework used in this study. The evaluation of the proposed methodology on the custom multimodality dataset, as well as the VGGSound dataset is provided in section 5. Moreover, in section 6 the paper showcases the real-time application of the proposed algorithm. Finally, a summary noting the advantages and disadvantages of the proposed method along with the scope for improvements and possible developments in the future are stated in the last section.
@@ 2.1 Continual Learning Strategies
Continual learning strategies can be broadly categorized into four major types: architectural, regularization, rehearsal, and hybrid strategies, each with it’s own pros and cons. Architectural strategy either employs a dynamic modular network where new layers gets added upon encounter of new tasks [10], [11] or performs parameter isolation to allocate a dedicated separate sub-space for each task within the same network [12], [13]. However, a major drawback of this strategy is the increase in the size of the architecture as new tasks are learned or the potential saturation of the architecture if a fixed model is used.
@@
Regularization strategies can be grouped into two: weight regularization and function regularization. Weight regularization constrains the update of the network weights by adding explicit regularization terms to the loss function, penalizing changes in each parameter based on it’s importance [14], [15]. Function regularization, on the other hand, employs distillation methods to maintain the knowledge between the various layers of the network. This typically involves treating the previously trained model as a teacher and the current model as a student, applying knowledge distillation (KD) techniques to align their outputs [16], [17]. However, regularizing the model can make it more difficult for the network to acquire new knowledge.
@@
Rehearsal [18], [19], & pseudo-rehearsal, [20], [21] approaches offer the most simple yet effective approach by storing a subset of the previously learned data in a memory buffer or by training a generative model to account for the previous information and replaying it during training on new tasks. However, storing the images in a replay buffer for rehearsal can raise privacy concerns. While using a generative model somewhat addresses the privacy issues, it may result in increased training time and computational demands.
@@
Hybrid strategies [22], [23] are a combination of any of the other three strategies to compensate for catastrophic forgetting. SynapNet [24] uses a triple model architecture along with a VAE-based pseudo-episodic memory for rehearsal and a sleep phase for memory re-organization. DualNet [25] uses a fast learner for supervised learning and a slow learner for unsupervised learning of task-agnostic general representation using semi-supervised learning.
@@
Non-exemplar CL strategies [8], [26] use the idea of replay-free sequential training, instead of using a modular network or regularizing the optimization function. These approaches store the prototypes of the past and new classes to compensate for catastrophic forgetting of the previously acquired knowledge. A pre-trained feature extractor provides the feature maps which are then used to calculate the mean and the covariance matrix for the respective classes.
@@ 2.2 Semi-Supervised Learning
SSL describes a class of algorithms that aim to learn from a combination of labeled and unlabeled data, which are typically assumed to be drawn from the same underlying distributions [27]. A central challenge in SSL lies in effectively leveraging both labeled and unlabeled data while mitigating the risks of overfitting, reducing the impact of noisy data, and ensuring meaningful improvements in model performance. Early advances in SSL using deep neural networks were driven by generative models such as autoencoders [28], variational autoencoders [29], and generative adversarial networks [30], which played a crucial role in extracting robust representations from partially labeled datasets.
Some SSL approaches [31], [32] focus on balancing two complementary loss terms: a supervised loss ($l_{s}$) applied to labeled samples and an unsupervised consistency regularization loss ($l_{ul}$) computed on the unlabeled data. These approaches are designed to ensure that the model learns not only from explicit supervision but also from consistency in predictions across augmented versions of unlabeled inputs. Two additional important approaches which have shown success in the context of deep learning are: self-training and conditional entropy minimization. Self-training, also known as pseudo-labeling [33], involves assigning labels to unlabeled data based on the predictions of a model initially trained on labeled examples. In contrast, conditional entropy minimization [34] encourages the model to make confident predictions on unlabeled samples by minimizing the entropy of their predicted class distributions. Although this does not ensure correctness due to the absence of true labels, it promotes decisiveness and reduces model uncertainty. In this work, we adopt the strategy of using cosine similarity between the feature representations of labeled and unlabeled data as a criterion for pseudo-labeling of the unlabeled samples. We refer to [6] for a comprehensive review on the topic.
@@ 2.3 Multimodality
Our experience of the world is inherently multimodal we perceive objects through vision, hear sounds, feel textures, smell odors, and taste flavors. To develop an AI agent with similar capabilities, several approaches have been explored [35], [36]. The simplest method to compensate for multiple domains is using the concept of joint representation where the unimodal signals are combined in the same representation space by simple concatenation [37]. [38] employed stacked denoising autoencoders to represent each modality individually and subsequently fusing them into a unified multimodal representation using an additional autoencoder layer.
Similarly, [39] suggested a multimodal autoencoder for semantic concept grounding. They incorporated a reconstruction loss for training the representation and added a term in the loss function to predict object labels using the representation.[40] used a deep belief network (DBN) for each modality and then combined them into joint representations for audio-visual emotion recognition. A similar architecture was adopted by [41] for audio-visual speech recognition (AVSR), [42] applied the same strategy for integrating audio and skeletal joint representations in gesture recognition. [43] recently utilized a multimodal CL dataset to learn complementary information from two different modalities and performed a weighted fusion of the feature maps obtained from the respective feature extractors. On the other hand, [44] developed an egocentric multimodal dataset using smart glasses equipped with sensors and utilized the temporal binding network (TBN) [45] to fuse information from multiple sensors into a unified feature vector for feature extraction and training.
However, a key limitation of these methods is their inability to incrementally adapt to new domains. Specifically, when training on multiple domains (such as sensor signals, visual images), the addition of a new modality often requires re-training the entire model from scratch. In contrast, our proposed exFeCAM algorithm addresses this challenge by treating each modality as a distinct domain, allowing for the incremental integration of new modalities without the need to restart training. Furthermore, our approach eliminates the need to store data in memory for rehearsal, effectively addressing privacy concerns and the storage constraints commonly associated with exemplar-based CL methods.
@@
Figure 1: Multimodal data comprising of two different domains, domain 1 contains the images of various objects in different orientations, lightning conditions, and different levels of occlusion by the operator and domain 2 contains the tactile sensor information for the corresponding objects. Moving from left to right the model trains in a class-incremental manner, while going from top-bottom it trains in a domain-incremental manner.